What is the fifth term of the sequence whose first term is a1 = 8 and whose nth term is an = 3an-1 + 7?
Answer by stanbon(75887)

What is the fifth term of the sequence whose first term is a1 = 8 and whose nth term is an = 3a(n-1) + 7?
------
a(5) = 3*a(4) + 7
------
a(1) = 8
a(2) = 3*a(2) + 7 = 24+7 = 31
a(3) = 3*a(2) + 7 = 93+7 = 100
a(4) = 3*a(3) + 7 = 300+7= 307
a(5) = 3*a(4) + 7 = 921+7 = 928
====================================
Cheers,
Stan H.